OpenStudy (anonymous):

Evaluate in^4(squrt)e

OpenStudy (anonymous):

\[\ln \sqrt[4]{e}\]

OpenStudy (anonymous):

AAAHHHHHH WASSUPPPP BROOOO

OpenStudy (anonymous):

waz up cuz

OpenStudy (anonymous):

uuuhhh ........

OpenStudy (anonymous):

wassup bro

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

cuh u aint king imprint

OpenStudy (anonymous):

?.?

OpenStudy (anonymous):

give me the answer choices

OpenStudy (anonymous):

\[\ln \sqrt[4]{e}\] there arent any

OpenStudy (anonymous):

A.e B.e^1/4 C.1/4 D.4

OpenStudy (anonymous):

so we just gotta figure it out

OpenStudy (anonymous):

so are the choices the teacher just gave us

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

A.e B.e^1/4 C.1/4 D.4

OpenStudy (anonymous):

oh easy b is the answer

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

same answer i think